OpenStudy (anonymous):

A department store sells a 4-pack of AAA batteries for $3.00 and an 8-pack of AAA batteries for $5.00. The cost of one battery from the 4-pack is what percent more than the cost of one battery from the 8-pack, before sales tax is applied?show work plz

OpenStudy (babalooo):

4pack = $3 $3/4, we get $0.75 for each battery 8pack = $5 $5/8 = $0.625 for each battery \[\frac{0.75 }{ 0.625 }= 1.2\] This shows that the cost of a 4pack battery is 120% of the cost of a battery from 8pack. To show the percent more in cost of a 4pack battery, simply subtract 100%. 120-100=20 So, a battery from 4pack is roughly 20% more expensive a battery from 8pack.